# Data-driven methods to discover stable linear models of the helicity injectors on HIT-SIU

### v1

@@ III Role of Data Driven Models
Data driven model discovery and control are growing in popularity in plasma physics. As plasma physics devices increase in complexity, data-driven system identification will be a critical tool for understanding and controlling the complex, nonlinear phenomena they exhibit. While deep learning methods such as deep reinforcement learning have been applied with great success in plasma physics [11, 46], these methods continue to struggle with interpretability, required training data, and uncertainty quantification [41]. Learning interpretable governing equations allows for comparisons with models derived from first-principles. In addition, these methods can be easily paired with statistical techniques to obtain average coefficient values, as well as the variation in coefficients across trials supporting qualification in licensing and other contexts. These methods are also lightweight and require relatively low compute time, making them prime candidates for real-time control. While there is some intuition about the nature of the evolution of the inductance on HIT-SIU, and previous iterations of the HIT-SI experiment, exact models for how the self and mutual inductance of the plasma evolves over time are not precisely known. However, there is a wealth of experimental data that can be used for the discovery of an improved linear model of the dynamics between the injector circuits and the plasma.
III.1 The Dynamic Mode Decomposition
@@
$\mathbf{x}_{k+1}=\mathbf{Ax}.$ (2)
Where $\mathbf{x}_{k}$ is an $n$-dimensional vector, where $n$ is the dimension of the state space, and $A\in\mathbb{R}^{n\times n}$. To find this matrix $A$, we arrange measurements, often referred to as snapshots, into two matrices $\bm{X}$ and $\bm{X}^{\prime}$. For $m$ measurements, $\bm{X}$,$\bm{X}^{\prime}$ $\in\mathbb{R}^{n\times m-1}$
$\mathbf{X}=\begin{bmatrix}|&|&&|\\ x_{1}&x_{2}&...&x_{m-1}\\ |&|&&|\end{bmatrix},\quad\mathbf{X^{\prime}}=\begin{bmatrix}|&|&&|\\ x_{2}&x_{3}&...&x_{m}\\ |&|&&|\end{bmatrix}.$ (3)
@@
The best-fit linear operator that satisfies the relationship in Eq. ( 4) is given as the solution to the matrix least squares problem
@@
Each eigen-pair corresponds to a distinct spatiotemporal mode, growing at a frequency $\mathrm{Re}(\lambda_{j})$ and oscillating at a frequency $\mathrm{Imag}(\lambda_{j})$. The states of the system can be reconstructed as a linear combination of the DMD modes
$\mathbf{x}_{k+1}=\sum_{i=1}^{r}b_{i}{\phi}_{i}\bm{\lambda}{{}_{i}}^{k}=\bm{\Phi B\Lambda}^{k},$ (11)
@@
$\mathbf{x}(t)=\sum_{i=1}^{r}b_{i}{\bm{\phi}}_{i}\exp(\omega_{i}t)=\bm{\Phi}\mathbf{B}\exp(\bm{\Omega}t),$ (12)
@@
III.2 The Optimized and Bagging Optimized DMD
While the exact DMD has shown great promise for evenly sampled and noise-free data, the algorithm often fails to correctly identify modes in noisy data. The presence of sensor noise obscures the true relationship between consecutive snapshot pairs, leading to incorrect forecasts and often unstable models of the dynamics. While there have been efforts to robustify the DMD to the effects of noise and actuation, such as forward-backward DMD [9], Total-Least-Squares DMD [20], Measure-preserving DMD [7], and DMD with control (DMDc) [40], we adopt the Optimized DMD (OPT-DMD) from Askham and Kutz [4] for performing DMD forecasting on noisy, experimental data as it offers the most robust framework to date for handling noise. Further, eigenvalue constraints for stable mode construction can be easily integrated into the framework. DMD, OPT-DMD and BOP-DMD algorithms are all included in the pyDMD package [24].
OPT-DMD reformulates the exact DMD as a nonlinear optimization problem and directly solves for the eigenvalues and eigenmodes of the DMD operator, denoted here as $\bm{\Omega}\in\mathbb{C}^{n\times n}$ and $\Phi_{B}=\bm{\Phi B}\in\mathbb{C}^{n\times m}$,
@@ III.2 The Optimized and Bagging Optimized DMD
This procedure also allows for the user to constrain the learned model to be linearly stable, and therefore avoid unstable forecasting. The physical injector circuits of HIT-SIU are linearly stable, so throughout the course of this work we enforce that the learned OPT/BOP DMD models are stable. Linear stability is enforced in the Matlab implementation of OPT-DMD during the optimization process by fitting an initial OPT-DMD model, checking if any of the learned eigenvalues have a positive real part, setting this real part to zero while preserving the complex component, and then using these new eigenvalues as an initial guess for another call of OPT-DMD.
OPT-DMD was further expanded upon by Sashidahar and Kutz [41] to include statistical bagging techniques (BOP-DMD), adding further robustness to OPT-DMD, and allowing for uncertainty quantification of the learned modes, and eigenvalues. As outlined in  [41], a BOP-DMD model is computed by fitting many OPT-DMD models to randomly selected subsets of the training data, and then taking the mean over the eigenvalues, eigenvectors, and weights of each of these models to arrive at an "average" model of the dynamics. Means are taken component wise for vectors, and are taken separately for the real and complex components of the eigenvalues and eigenvectors. To ensure that eigenpairs that are trying to capture the same underlying physics are averaged with one another, the eigenpairs are sorted during each iteration of BOP-DMD. The authors of this work experimented with using medians as well, but found this did not increase the performance of BOP-DMD for this particular dataset.
IV Results
@@
After extensive testing, we have found that including nine DMD modes for a plasma model, and five DMD modes for a vacuum model, provide the best models for reconstruction of the training shot, and the prediction of future shots. We also enforce that the BOP-DMD model is linearly stable. This is achieved in the OPT-DMD code in  [4] by examining the real parts of the DMD eigenvalues at each step in the optimization, and setting any positive real parts to zero. The complex part of the eigenvalues are left intact, and then these new eigenvalues are used as an initial guess for the next step in the optimization.
IV.1 BOP and OPT DMD on Simulated data
@@
Figure 13: OPT-DMD is trained on the last quarter of a simulated vacuum shot once the power supplies have turned off, and incorrectly learns the dynamics matrix that generated the waveform.
IV.2 BOP-DMD on Vacuum Shots
@@
When beginning to examine the methodology outlined above, we initially tested the capabilities of BOP-DMD on data that was known to be generated by a linear dynamical system: the flux circuits in vacuum. Coupled LC circuits are a classical linear dynamical system, and thus present a test-bed for the performance of BOP-DMD under real-world experimental conditions. The results of BOP-DMD trained on this kind of data can be seen in Figs. 14 and 15. In Fig. 14, we show the performance of BOP-DMD trained on the last millisecond of a vacuum shot as the power supplies turn off. The model trained in Fig. 14 is then evaluated on the previous vacuum shot from that day Fig. 15. These results indicate that BOP-DMD is able to learn the underlying dynamics of the vacuum circuits, while not being subject to over fitting. Further, BOP-DMD is able to learn the correct structure of an underlying operator, and reduce the impact of spurious entries through statistical bagging.
@@
Figure 14: BOP-DMD trained on the fifth shot of August 16, 2022 in vacuum. This model was trained on the last millisecond of a shot once the power supplies had turned off. The model shown used five DMD modes (much less than the full 12 modes of the analytic model), and used 20 bags.
@@
Figure 15: This is the same model as used in Fig.  14, but is now simulated using a power supply waveform from the previous shot in vacuum (the fourth shot of August 16th, 2022). This model generalizes from shot to shot.
IV.3 BOP-DMD With a Plasma Present
@@
We now evaluate our methodology on shots where a plasma is present. By augmenting the states of our system to also include the toroidal plasma current, we train a BOP-DMD model on the last millisecond of a plasma shot. During this final millisecond of the shot, the toroidal plasma current decays from it’s steady state behavior as the plasma in the confinement volume begins to dissipate. Again, due to the limitations of BOP-DMD, this methodology is unable to access the steady state interaction between the flux circuits and the spheromak. However, this method still provides reasonable results; Fig. 16 shows an example of a BOP-DMD model fit on the last millisecond of a plasma shot and then evaluated on the rest of the shot. Further, this model is able to be extrapolated to other plasma shots taken on the same day Fig. 17.However, there error between the predicted current and the measured current in the third flux coil increases between the training shot, and the test shot. Because of the small differences in the relative phasing of the power supplies between the two shots, there is a different mutual inductance between the circuits and the plasma. As previously mentioned, one of the primary limitations of the BOP-DMD method is the inability to disambiguate the effects of actuation on the dynamics of a system. This leads to a fundamental gap in the training data available to the method for capturing the coupling between the injectors circuits and the plasma. Future work should focus on not only incorporating actuation into BOP-DMD, but also stringing together multiple shots for a wider array of training data. We do notice however that the operator learned by BOP-DMD does not resemble the same sparse structure that is present in the analytic plasma model Fig. 18. Despite losing the sparse structure of the dynamics matrix, BOP-DMD does identify reasonable dominant modes for the injector circuits. From Fig.  19, one can see that both the analytic model, and the BOP-DMD model identify dominant modes are roughly 20 kHz, which matches the resonant frequency of the injector circuits. It is also useful to examine the different predictions provided by each individual OPT-DMD model in the 20 iterations of BOP-DMD  20.
@@
Figure 16: BOP-DMD trained on the last millisecond of the 9th shot of August 16th, 2022. This is a shot when a plasma is present. Even though BOP-DMD is trained on the final transient at the end of the shot when the spheromak is decaying, the model is able to capture much of the flux circuit dynamics in the beginning of the shot. This model utilizes nine DMD modes. To the right of the vertical blue line, there is no plasma present, and to the right, a plasma is present.
@@
Figure 17: BOP-DMD model trained on last millisecond of shot 220816009 simulated with a new power supply waveform from the previous shot and evaluated against the ground truth. Even though the BOP-DMD model had no training data from this shot, the model still captures the flux circuit dynamics. To the right of the vertical blue line, there is no plasma present, and to the right, a plasma is present.
@@
Figure 18: Comparison of the structure of BOP-DMD $\bm{A}_{p}$ matrix and analytic plasma model.
@@
IV.4 Implementation of BOP-DMD in a Control Loop
Now that we have demonstrated that BOP-DMD can learn linear models of the flux circuits both in the vacuum case and the plasma case, we will demonstrate the effectiveness of this model as a plant in an LQG feedback loop. Because this model is of reduced rank (9 DMD modes, only five of which are orthogonal), we begin with taking an SVD of the BOP-DMD matrix, and projecting onto the first five columns of $\mathbf{U}$,
@@ IV.4 Implementation of BOP-DMD in a Control Loop
V Conclusion
In this work, we have demonstrated the advantages of linear control for the flux and voltage circuits on the HIT-SIU device, and the role BOP-DMD plays in discovering linear models of the complex interaction between the flux and voltage circuits and the spheromak. While BOP-DMD has limitations in terms of disambiguating the affects of actuation on data, there are still many advantages to using BOP-DMD over first-principles linear models. BOP-DMD does not require knowledge of often hard-to-measure system parameters (in this case the inductance, mutual inductance, and capacitance of the various circuit components), and BOP-DMD is able to be tuned to shot-to-shot experimental conditions with ease. BOP-DMD’s ability to provide interpretable, linear models of often nonlinear circuit behavior with uncertainty quantification make it a powerful tool for use in linear optimal control and estimation.
We have also demonstrated in this work the advantages of linear optimal control to model-agnostic control schemes such as PID control, on the HIT-SIU device. By pairing linear optimal control, with fully data-driven system identification, we have demonstrated that a model-based control scheme can accurately model the flux circuit dynamics, and control the behavior of the model in a simulation environment. Despite the advancements put forth in this work, the authors acknowledge room for improvement in this methodology, specifically with regards to the need for DMD variants that are able to robustly remove the affect of actuation on a dynamical system in the presence of noise. This work highlights the need for new approaches such as a combination of BOP-DMD and DMDc, that leverages reformulating the classical DMD problem as a best fit of complex exponential’s, while at the same time being able to remove the affect of a known control signal from the data.
VI Acknowledgments

### v3

@@ IV Role of Data-Driven Models
Data-driven model discovery and control are growing in popularity in plasma physics. As plasma physics devices increase in complexity, data-driven system identification will be a critical tool for understanding and controlling the complex, nonlinear phenomena they exhibit. While deep learning methods such as deep reinforcement learning have been applied with great success in plasma physics [12, 50], these methods continue to struggle with interpretability, required training data, and uncertainty quantification [47]. Learning interpretable governing equations allows for comparisons with models derived from first-principles. In addition, these methods can be easily paired with statistical techniques to obtain average coefficient values, as well as the variation in coefficients across trials supporting qualification in licensing and other contexts. These methods are also lightweight and require relatively low compute time, making them prime candidates for real-time control. While there is some intuition about the nature of the evolution of the inductance on HIT-SIU, and previous iterations of the HIT-SI experiment, exact models for how the self and mutual inductance of the plasma evolve over time are not precisely known. However, there is a wealth of experimental data that can be used for the discovery of an improved linear model of the dynamics between the injector circuits and the plasma.
IV.1 The Dynamic Mode Decomposition
@@
$\mathbf{x}_{k+1}=\mathbf{Ax}_{k}.$ (3)
Where $\mathbf{x}_{k}$ is an $n$-dimensional vector, where $n$ is the dimension of the state space, and $A\in\mathbb{R}^{n\times n}$. To find this matrix $A$, we arrange measurements, often referred to as snapshots, into two matrices $\bm{X}$ and $\bm{X}^{\prime}$. For $m$ measurements, $\bm{X}$$\bm{X}^{\prime}$ $\in\mathbb{R}^{n\times m}$
$\mathbf{X}=\begin{bmatrix}|&|&&|\\ x_{0}&x_{1}&...&x_{m-1}\\ |&|&&|\end{bmatrix},\quad\mathbf{X^{\prime}}=\begin{bmatrix}|&|&&|\\ x_{1}&x_{2}&...&x_{m}\\ |&|&&|\end{bmatrix}$. (4)
@@
The best-fit linear operator that satisfies the relationship in Eq. ( 5) is given as the solution to the matrix least squares problem
@@
Each eigen-pair corresponds to a distinct spatiotemporal mode, growing or damping at a rate $\mathrm{Re}(\lambda_{j})$, and oscillating at a frequency $\mathrm{Imag}(\lambda_{j})$. The states of the system can be approximated as a linear combination of the DMD modes
$\mathbf{x}_{k+1}\approx\sum_{i=1}^{r}b_{i}{\phi}_{i}\bm{\lambda}{{}_{i}}^{k}=\bm{\Phi}\mathbf{B}\bm{\Lambda}^{k},$ (12)
@@
$\mathbf{x}(t)\approx\sum_{i=1}^{r}b_{i}{\bm{\phi}}_{i}\exp(\omega_{i}t)=\bm{\Phi}\mathbf{B}\exp(\bm{\Omega}t),$ (13)
@@
IV.2 The Optimized and Bagging Optimized DMD
While the exact DMD has shown great promise for evenly sampled and noise-free data, the algorithm often fails to correctly identify modes in noisy data. The presence of sensor noise obscures the true relationship between consecutive snapshot pairs, leading to incorrect forecasts and often unstable models of the dynamics. There have been many efforts to improve upon exact DMD with respect to the effects of noise and actuation, such as forward-backward DMD [10], Total-Least-Squares DMD [21], Measure-preserving DMD [8], Higher Order DMD [35] and DMD with control (DMDc) [44]. We adopt the Optimized DMD (OPT-DMD) from Askham and Kutz [5] for performing DMD forecasting on noisy, experimental data as it offers the most robust framework to date for handling noise. Further, eigenvalue constraints for stable mode construction can be easily integrated into the framework. DMD, OPT-DMD and BOP-DMD algorithms are all included in the pyDMD package [26].
OPT-DMD reformulates the exact DMD as a nonlinear optimization problem and directly solves for the eigenvalues and eigenmodes of the DMD operator, denoted here as $\bm{\Omega}\in\mathbb{C}^{n\times n}$ and $\bm{\Phi_{B}}=\bm{\Phi B}\in\mathbb{C}^{n\times m}$,
@@ IV.2 The Optimized and Bagging Optimized DMD
This procedure also allows for the user to constrain the learned model to be linearly stable, and therefore avoid unstable forecasting. The physical injector circuits of HIT-SIU are linearly stable, so throughout the course of this work we enforce that the learned OPT/BOP DMD models are stable. Linear stability is enforced in the Matlab implementation of OPT-DMD that was developed for  [5]. The optimizer fits an initial OPT-DMD model, and checks if any of the learned eigenvalues have a positive real part. If so, the positive real components of the eigenvalues are set to zero, while preserving the complex component. These new eigenvalues are then used as an initial guess for another call of OPT-DMD.
OPT-DMD was further expanded upon by Sashidhar and Kutz [47] to include statistical bagging techniques (BOP-DMD), adding further robustness to OPT-DMD, and allowing for uncertainty quantification of the learned modes, and eigenvalues. As outlined in  [47], a BOP-DMD model is computed by fitting many OPT-DMD models to randomly selected subsets of the training data, and then taking the mean over the eigenvalues, eigenvectors, and weights of each of these models to arrive at an "average" model of the dynamics. Means are taken component wise for vectors, and are taken separately for the real and complex components of the eigenvalues and eigenvectors. To ensure that eigenpairs that are trying to capture the same underlying physics are averaged with one another, the eigenpairs are sorted during each iteration of BOP-DMD. The authors of this work experimented with using medians as well, but found this did not increase the performance of BOP-DMD for this particular dataset.
V Results
@@
After extensive testing, we have found that including nine DMD modes for a plasma model, and five DMD modes for a vacuum model, provide the best models for reconstruction of the training shot, and the prediction of future shots.
V.1 BOP and OPT DMD on Simulated data
@@
Figure 13: OPT-DMD is trained on the last quarter of a simulated vacuum shot once the power supplies have turned off. There are spurious entries to the DMD operator computed by the OPT-DMD that are able to be rectified with the addition of statistical bagging.
V.2 BOP-DMD on Vacuum Shots
@@
When beginning to examine the methodology outlined above, we initially tested the capabilities of BOP-DMD on data that was known to be generated by a linear dynamical system: the flux circuits in vacuum. Coupled LC circuits are a classical linear dynamical system, and thus present a test-bed for the performance of BOP-DMD under real-world experimental conditions. The results of BOP-DMD trained on this kind of data can be seen in Figs. 14 and 15. In Fig. 14, we show the performance of BOP-DMD trained on the last millisecond of a vacuum shot as the power supplies turn off, and then evaluated on the entire shot by feeding in full power supply waveform. The model trained in Fig. 14 is then evaluated on the previous vacuum shot from that day Fig. 15. These results indicate that BOP-DMD is able to learn the underlying dynamics of the vacuum circuits, while not being subject to over fitting. Further, BOP-DMD is able to learn the correct structure of an underlying operator, and reduce the impact of spurious entries through statistical bagging.
@@
Figure 14: BOP-DMD trained on the fifth shot of August 16, 2022 in vacuum. This model was trained on the last millisecond of the shot once the power supplies had turned off, and then evaluated by simulating the model with the full power supply waveform comparing to the original shot. The model shown used five DMD modes (much less than the full 12 modes of the analytic model), and used 20 bags.
@@
Figure 15: This is the same model as used in Fig.  14, but is now simulated using a power supply waveform from the previous shot in vacuum (the fourth shot of August 16th, 2022). This model generalizes from shot to shot from the same day.
V.3 BOP-DMD With a Plasma Present
@@
We now evaluate our methodology on shots where a plasma is present. By augmenting the states of our system to also include the toroidal plasma current, we train a BOP-DMD model on the last millisecond of a plasma shot. During this final millisecond of the shot, the toroidal plasma current decays from it’s steady state behavior as the plasma in the confinement volume begins to dissipate. Again, due to the limitations of BOP-DMD, this methodology is unable to access the steady state interaction between the flux circuits and the spheromak. However, this method still provides reasonable results; Fig. 16 shows an example of a BOP-DMD model fit on the last millisecond of a plasma shot and then evaluated on the rest of the shot. Further, this model is able to be applied to other plasma shots taken on the same day Fig. 17. However, the error between the predicted current and the measured current in the third flux coil increases between the training shot, and the test shot. Because of the small differences in the relative phasing of the power supplies between the two shots, there is a different mutual inductance between the circuits and the plasma. As previously mentioned, one of the primary limitations of the BOP-DMD method is the inability to disambiguate the effects of actuation on the dynamics of a system. This leads to a fundamental gap in the training data available to the method for capturing the coupling between the injectors circuits and the plasma. Future work should focus on not only incorporating actuation into BOP-DMD, but also stringing together multiple shots for a wider array of training data. We do notice however that the operator learned by BOP-DMD does not resemble the same sparse structure that is present in the analytic plasma model Fig. 18. Despite losing the sparse structure of the dynamics matrix, the modes of the BOP-DMD model oscillate at similar frequencies to the analytic model, despite decaying faster in general. From Fig.  19, one can see that both the analytic model, and the BOP-DMD model identify dominant modes at roughly 20 kHz, which matches the resonant frequency of the injector circuits. It is also useful to examine the different predictions provided by each individual OPT-DMD model in the 20 iterations of BOP-DMD  20.
@@
Figure 16: BOP-DMD trained on the last millisecond of the 9th shot of August 16th, 2022. This is a shot when a plasma is present. Even though BOP-DMD is trained on the final transient at the end of the shot when the spheromak is decaying, the model is able to capture much of the flux circuit dynamics in the beginning of the shot. This model utilizes nine DMD modes. To the left of the vertical blue line ($\approx.6$ ms), there is no plasma present, and to the right, a plasma is present.
@@
Figure 17: BOP-DMD model trained on last millisecond of shot 220816009 simulated with a new power supply waveform from the previous shot and evaluated against the ground truth. Even though the BOP-DMD model had no training data from this shot, the model still captures the flux circuit dynamics. To the left of the vertical blue line ($\approx.6$ ms), there is no plasma present, and to the right, a plasma is present.
@@
Figure 18: Comparison of the structure of BOP-DMD $\mathbf{A}_{p}$ matrix and analytic plasma model.
@@
V.4 Implementation of BOP-DMD in a Control Loop
Now that we have evaluated the ability of BOP-DMD to learn linear models of the flux circuits both in the vacuum case and the plasma case, we will evaluate the effectiveness of this model as a plant in an LQG feedback loop. Because this model is of reduced rank , we begin with taking an SVD of the BOP-DMD matrix, and projecting onto the first five columns of $\mathbf{U}$,
@@ V.4 Implementation of BOP-DMD in a Control Loop
VI Conclusion
In this work, we have demonstrated the advantages of linear control for the flux and voltage circuits on the HIT-SIU device, and the role BOP-DMD plays in discovering linear models of the complex interaction between the flux and voltage circuits and the spheromak. While BOP-DMD has limitations in terms of disambiguating the effects of actuation on data, there are still many advantages to using BOP-DMD over first-principles linear models. BOP-DMD does not require knowledge of often hard-to-measure system parameters (in this case the inductance, mutual inductance, and capacitance of the various circuit components), and BOP-DMD is able to be tuned to shot-to-shot experimental conditions with ease. BOP-DMD’s ability to provide interpretable, linear models of often nonlinear circuit behavior with uncertainty quantification make it a powerful tool for use in linear optimal control and estimation.
We have also demonstrated in this work the advantages of linear optimal control to model-agnostic control schemes such as PID control, on the HIT-SIU device. By pairing linear optimal control, with fully data-driven system identification, we have demonstrated that a model-based control scheme can accurately model the flux circuit dynamics, and control the behavior of the model in a simulation environment. Despite the advancements put forth in this work, the authors acknowledge room for improvement in this methodology, specifically with regards to the need for DMD variants that are able to robustly remove the affect of actuation on a dynamical system in the presence of noise. This work highlights the need for new approaches such as a combination of BOP-DMD and DMDc, that leverages reformulating the classical DMD problem as a best fit of complex exponentials, while at the same time being able to remove the affect of a known control signal from the data.
VII Acknowledgments